# A Survey on Data Selection for Language Models

### v2

@@ 7 Data Selection for Task-specific Fine-tuning
There have been a number of works that utilize auxiliary data, data from a source other than the target task, to improve a model’s ability to generalize to the target distribution. In these settings, the goal of data selection is to determine the optimal auxiliary data that improves a model’s performance on the target distribution, and because we have access to data from the target distribution the utility function for these methods is often a form of similarity between the auxiliary and target data. These methods tend to have some similarities with domain-specific data selection for pretraining data (§3.4) with the main difference being that in a fine-tuning setting, more computationally expensive utility functions are feasible.
@@
While Phang et al. (2018) and Pruksachatkun et al. (2020) focus on task transfer from the auxiliary to target data, Iter & Grangier (2021) and Grangier & Iter (2022) focus on domain adaptation from the auxiliary to target data. Iter & Grangier (2021) find that the best data selection method for domain adapatation uses a domain classifier as the utility function, a model that is trained to discriminate between data from the target distribution and from a general distribution, similar to methods on domain-specific selection for pretraining (§3.4). Grangier & Iter (2022) further explore the use of three separate data selection methods: importance sampling, contrastive data selection, and influence functions, finding that all three methods select for very similar data points.
@@
More recently, methods that select from auxiliary data train on both the auxiliary and target data simultaneously. For example, FETA (Albalak et al., 2022b) performs a comprehensive task transfer study similar to Pruksachatkun et al. (2020), training a model on all auxiliary-target pairs, and finding that similarity between the auxiliary and target task’s label-space (e.g. binary classification, multi-class classification, etc.) can be used as a good heuristic utility function. However, using the label-space is only possible when the structure of the target task is very clear, and is only useful when the label-space is fairly unique. Additionally, the structure of a task does not fully describe the data, and this method is not feasible at large scales as it requires human judgment. DEFT-Few (Ivison et al., 2023a) improves on the ideas from FETA by using a model-based utility function to select the 500 auxiliary data points that are nearest neighbors to the target dataset. Skill-it (Chen et al., 2023c) also builds on the ideas from FETA (Albalak et al., 2022b), defining tasks as different skills that a model can learn, and learning a pairwise correlation between skills by training on each pair of skills to find when skill A benefits skill B. Based on this, they create a graph of skills ordered by dependencies and define a data mixing algorithm that incorporates the loss on the target task. However, training a separate model on every pairwise combination of tasks may be very inefficient. Albalak et al. (2023b) improves upon the inefficiencies of Skill-it by formulating the auxiliary data selection problem as a multi-armed bandit, where the utility function is an aggregated reward function of the cosine similarity and magnitude similarity between gradients of batches from the auxiliary and target datasets. By formulating the auxiliary data selection problem as a multi-armed bandit, their algorithm runs online without requiring any knowledge of the auxiliary data prior to training. While Albalak et al. (2023a) design a gradient-based utility function for data mixing, Xia et al. (2024b) use a similar gradient-based utility to assign values to individual auxiliary data points. To improve the efficiency of assigning utilities to each data point, they only calculate gradients for the parameter-efficient LoRA (Hu et al., 2022) modules and further reduce the gradient dimension through a random projection.
@@
Another setting where data selection can be applied beneficially for task-specific fine-tuning is when we wish to select a subset of the target dataset to improve data efficiency. Methods in this category have a similar motivation to fuzzy data deduplication from §3.5, but because the target datasets when performing fine-tuning are significantly smaller than pretraining corpora, the methods used here can apply more complex, computationally expressive utility functions.
Maharana et al. (2024) propose a method of diversifying data based on the expected difficulty. They represent the dataset as a graph, where individual nodes (data points) are connected to their k-nearest neighbors in representation space and use variability (Swayamdipta et al., 2020) as the utility function for each individual data point. Next, they perform a forward message passing process which propagates information on the utility function of a data point to it’s neighbors. Finally, they perform data selection through an iterative reverse message passing process, where at each step the highest scoring data point is kept, and all it’s neighbors get down-weighted to encourage the selection of diverse data points. Lang et al. (2022) also use a graph representation as part of their utility function, determining which weakly-labeled data points to keep based on their nearest neighbors, and the label of their nearest neighbors. They only keep data points whose nearest neighbors have mostly the same label, thereby using the representations geometry to identify accurate subsets and removing ambiguous data points. Bejan et al. (2023) develop the task-agnostic self-influence score as a utility function for filtering data (and designing a curriculum). The intuition behind the self-influence score is: if not training on data point $x^{(i)}$ leads to a significantly higher loss on that data point, then there must not be many supporting data points around $x^{(i)}$ and it is likely to be an outlier, and therefore a low-quality data point. They suggest that self-influence functions can identify data with label noise, out-of-distribution data, data that is ambiguous, and difficult to learn samples, and removing such data can improve the models performance.
Many methods of improving data efficiency exist in general machine learning settings (e.g. data pruning and coreset selection), but most have exclusively been developed and evaluated in the computer vision domain. Azeemi et al. (2023) is one of the limited tests of data pruning on language tasks, where they observe that selecting for difficult to learn examples leads to improved model performance, while reducing the dataset size by up to 60%. The many other works on data pruning and coreset selection (further discussed in §9) may provide effective methods of improving data efficiency for task-specific fine-tuning, but could require some alterations to work for NLP tasks.
@@
Finally, an important problem that data selection can solve is to improve a models robustness by ensuring that a model is not biased towards any particular subpopulation of data and to avoid relying on spurious correlations. Because these methods concern themselves with subgroups of data, they share similarities with data mixing (§3.8). Sagawa et al. (2020) introduce group distributionally robust optimization (group DRO) to train models that minimize the worst-case loss over groups in the training data. As their utility function, they use a notion of the generalization gap (the difference between training and validation loss) and prioritize training on data from those subgroups which have a large generalization gap. Haghtalab et al. (2022) also aim to minimize the worst-case error of a model across subgroups, while also aiming to improve the data efficiency by minimizing the number of samples required for each subgroup. They formulate the multi-distribution learning problem as a zero-sum game and aim to learn the minimax equilibrium by introducing an online learning algorithm, Resampling-based Multi-Distribution Learning (R-MDL) that uses estimates of the training error and generalization error as the utility.
@@
Compared with data selection for pretraining, the goals of data selection for task-specific fine-tuning are very similar: improving data efficiency, selecting a diverse set of samples, and reducing a models bias towards any particular subgroup. However, the task-specific datasets used in fine-tuning are often small enough that methods with high computational requirements are plausible. Though many of the methods presented in this section may be beneficial for pretraining (or other settings), they may be unreasonably costly to use on web-scale data. For example, Azeemi et al. (2023) requires training a model on the full dataset prior to determining which data points to keep/remove, and Haghtalab et al. (2022) assumes that the model will be validated on each subgroup at every step of the learning algorithm. Although these ideas may be computationally prohibitive to be used as-is, it may be possible to adjust them to work at larger scales by reducing the frequency of updates, or using smaller models as proxies.
@@ 8 Data Selection for Other Domains
This survey focuses on data selection for language models, but data selection is an active area of research in computer vision, vision-and-language and broadly in machine learning. Some of the data selection methods covered by this survey are domain-specific (e.g. the total number of characters in $x^{(i)}$), while others are domain agnostic (e.g. selecting data most similar to a target distribution). Therefore, some of the methods previously presented in this work may be more applicable in other domains, and vice-versa. Similar to language models, data selection methods have been applied at multiple stages of model training in other domains (e.g. vision-language model pretraining) and for similar purposes (e.g. data efficiency, model performance, and reducing bias). In this section we will briefly discuss some of the prominent uses of data selection in non-language domains, and compare and contrast them with the previously discussed methods.
@@
In addition to data collection and filtering, data mixing is also a topic of concern for non-language pretraining. Piergiovanni et al. (2023) propose an online data mixing method for vision-language models that mixes a variety of tasks together (e.g. image captioning or visual question answering). Their method, dynamic difficulty sampling, calculates the loss of each task $t$, $\mathcal{L}_{t}$, and weights each task proportionally with the total loss $L$, getting the final sampling weight as $S_{t}=\dfrac{\mathcal{L}_{t}}{L}$. This mixing method maintains some similarities to the online data mixing methods from language model pretraining (Xia et al., 2024a; Albalak et al., 2023a; Schioppa et al., 2023).
@@
The dominant paradigm in machine learning’s history has been task-specific fine-tuning, and many methods of data selection have been proposed for this setting. For the most part, the methods of data selection for non-language domains are quite similar to the methods for language. Generally speaking, these methods use the output of a model as the representation upon which the utility function is built. In an early work on support vector machines, Wang et al. (2005) devise a method of selecting data that reduces the training set size while maintaining generalization by filtering out data points which do not affect the final decision function. This idea resonates in the language domain, with similarities to methods that use loss (Albalak et al., 2023a) or methods that attempt to estimate data difficulty (Mindermann et al., 2022). A more recent direction has been how to best utilize synthetic data, and Yuan et al. (2024a) demonstrate that generating large quantities of synthetic images and selecting data using distribution matching (matching the synthetic data to the target data using maximum mean discrepancy) leads to models that perform well for image classification with synthetic data only, as well as when mixing real and synthetic data.
@@
A number of methods have been developed to improve the data selection on the level of individual batches, sometimes called data difficulty estimation. One of the motivations of these methods is that they use the current model’s state to determine which data to select, but this can add significant computational overhead. Additionally, because these methods are run alongside the model training, their cost is not amortized across training runs as it would be for static datasets. Chang et al. (2017) propose two methods for Active Bias that emphasize training on data points where the model is uncertain, based on their prediction variance or closeness to the decision threshold. Similarly, Song et al. (2020) propose Recency Bias, a method that further emphasizes training on uncertain data points that were predicted inconsistently in recent iterations of training. Finally, (Mindermann et al., 2022) present Reducible Holdout Loss Selection (RHO-Loss), a method of adaptive batch selection that, intuitively, avoids selecting data points that are redundant, noisy, or less relevant. RHO-Loss was validated on a very small sample of NLP tasks, task-specific fine-tuning on SST2 (Socher et al., 2013) and CoLA (Warstadt et al., 2019), but has not been demonstrated to work in other learning settings, and Kaddour et al. (2024) show that it is not computationally efficient.
Adaptive batch selection methods are similar to online data selection, except that they are often performed on the level of individual data points, rather than groups, domains, or tasks. This particular direction has found minimal success with LLMs, in particular for pretraining, but the success of such methods in other domains suggests that it would be worthwhile adapting these methods to language.
@@
There have been a number of recent works that aim to prepare fair datasets. Biswas & Rajan (2021) analyze stage-specific fairness (SF), measuring fairness metrics with and without certain data preprocessing stages, and provided experimental results. González-Zelaya et al. (2023) introduced FairPipes, a heuristic approach to find an optimal data preparation pipeline configuration. A number of works have also developed software tools for tracing distribution distortions linked to fairness in the data preparation workflow (Yang et al., 2020a; Grafberger et al., 2021; Schelter et al., 2019).
Additionally, there exist fairness-aware balancing techniques (Yan et al., 2020; Sonoda, 2023; Chakraborty et al., 2021; Salazar et al., 2021; Yang et al., 2020b) as well as reweighting/resampling methods aimed at ensuring fairness (Kamiran & Calders, 2012; Jiang & Nachum, 2020). In NLP, using balanced training for fairness has been explored (Wang et al., 2019; Han et al., 2022). Jeong et al. (2022b) showed that different proportions of demographics can lead to nontrivial differences in fairness. To alleviate the challenge of gathering additional data, recent research has focused on designing adversarial networks. These networks are developed to acquire a balanced representation, which helps in minimizing the adverse effects of sensitive attributes (Madras et al., 2018; Adel et al., 2019).
There are a number of works presented in this survey that discuss filtering and cleaning data, but another common issue when collecting new data is missing values. When a data entry contains missing values, should the entry be dropped, or can the missing value be imputed? Recent and concurrent research papers that study fairness in missing values (Caton et al., 2022; Schelter et al., 2019; Fernando et al., 2021; Mansoor et al., 2022; Jeanselme et al., 2022; Zhang & Long, 2021), which primarily focus on empirically assessing how dropping or various imputation algorithms affect fairness in downstream task. Jeong et al. (2022a) theoretically showed how discrimination risks can arise during data imputation and developed an algorithm called Fair MIP Forest that integrates missing value treatment and learning into a single process, and impose a fairness regularizer in an end-to-end process.
@@
Recently, the importance of data selection and quality has gained notoriety, and thus a few efforts have been developed to try and lower the barrier to entry for research to be done. One such effort is DataComp (Gadre et al., 2023), a benchmark for multimodal dataset design. DataComp provides two tracks for participation: the filtering track, and the bring-your-own-data track. The filtering track is of particular interest as it gives participants a common dataset (12.8B image-text pairs from Common Crawl), and the participants goal is to determine the best subset to train a model on. Additionally, Gadre et al. (2023) provide over 300 baseline experiments, and find that, similar to the language-only setting, a smaller, more selective dataset can lead to models that generalize better than larger datasets. Of course, this still comes with the same limitations as all previously discussed sections, where the exact evaluation setting plays a large role in what data is “best”. Specifically, DataComp evaluates models’ abilities to perform image classification and retrieval tasks, but does not evaluate generative capabilities. This suggests that distribution matching methods may fare well on DataComp, while an evaluation of generative capabilities may prefer data selection methods that favor distribution diversification. DataPerf (Mazumder et al., 2023) provides another entry for data selection research. The DataPerf benchmark contains 4 data selection settings: vision, speech, debugging, and language. The various settings allow researchers to test their algorithms for selection, augmentation, quality assesment, and cleaning. Mazumder et al. (2023) find that automated methods tended to perform better than manual cleaning or filtering, specifically, they found that submissions succeeded with automated methods for recognizing noisy images and labels, identifying mislabeled images, correcting class imbalance, and selecting and enhancing images from the long-tailed distribution of classes.

### v3

@@ 7 Data Selection for Task-specific Fine-tuning
There have been a number of works that utilize auxiliary data (i.e. data from a source other than the target task) to improve a model’s ability to generalize to the target distribution. In these settings, the goal of data selection is to determine the optimal auxiliary data that improves a model’s performance on the target distribution, and because we have access to data from the target distribution the utility function for these methods is often a form of similarity between the auxiliary and target data. These methods tend to have some similarities with domain-specific data selection for pretraining data (§3.4) with the main difference being that in a fine-tuning setting, more computationally expensive utility functions are feasible.
@@
While Phang et al. (2018) and Pruksachatkun et al. (2020) focus on task transfer from the auxiliary to target data, Iter & Grangier (2021) and Grangier & Iter (2022) focus on domain adaptation from the auxiliary to target data. Iter & Grangier (2021) find that the best data selection method for domain adaptation uses a domain classifier as the utility function, a model that is trained to discriminate between data from the target distribution and from a general distribution, similar to methods on domain-specific selection for pretraining (§3.4). Grangier & Iter (2022) further explore the use of three separate data selection methods: importance sampling, contrastive data selection, and influence functions, finding that all three methods select for very similar data points.
@@
More recently, methods that select from auxiliary data train on both the auxiliary and target data simultaneously. For example, FETA (Albalak et al., 2022b) performs a comprehensive task transfer study similar to Pruksachatkun et al. (2020), training a model on all auxiliary-target pairs, and finding that similarity between the auxiliary and target task’s label-space (e.g., binary classification, multi-class classification, etc.) can be used as a good heuristic utility function. However, using the label-space is only possible when the structure of the target task is very clear, and is only useful when the label-space is fairly unique. Additionally, the structure of a task does not fully describe the data, and this method is not feasible at large scales as it requires human judgment. DEFT-Few (Ivison et al., 2023a) improves on the ideas from FETA by using a model-based utility function to select the 500 auxiliary data points that are nearest neighbors to the target dataset. Skill-it (Chen et al., 2023c) also builds on the ideas from FETA (Albalak et al., 2022b), defining tasks as different skills that a model can learn, and learning a pairwise correlation between skills by training on each pair of skills to find when skill A benefits skill B. Based on this, they create a graph of skills ordered by dependencies and define a data mixing algorithm that incorporates the loss on the target task. However, training a separate model on every pairwise combination of tasks may be very inefficient. Albalak et al. (2023b) improves upon the inefficiencies of Skill-it by formulating the auxiliary data selection problem as a multi-armed bandit, where the utility function is an aggregated reward function of the cosine similarity and magnitude similarity between gradients of batches from the auxiliary and target datasets. By formulating the auxiliary data selection problem as a multi-armed bandit, their algorithm runs online without requiring any knowledge of the auxiliary data prior to training. While Albalak et al. (2023a) design a gradient-based utility function for data mixing, Xia et al. (2024b) use a similar gradient-based utility to assign values to individual auxiliary data points. To improve the efficiency of assigning utilities to each data point, they only calculate gradients for the parameter-efficient LoRA (Hu et al., 2022) modules and further reduce the gradient dimension through a random projection.
@@
Another setting where data selection can be applied beneficially for task-specific fine-tuning is when we wish to select a subset of the target dataset to improve data efficiency. Methods in this category have a similar motivation to fuzzy data deduplication from §3.5, but because the target datasets when performing fine-tuning are significantly smaller than pretraining corpora, the methods used here can apply more complex, computationally expensive utility functions.
Maharana et al. (2024) propose a method of diversifying data based on the expected difficulty. They represent the dataset as a graph, where individual nodes (data points) are connected to their k-nearest neighbors in representation space and use variability (Swayamdipta et al., 2020) as the utility function for each individual data point. Next, they perform a forward message passing process which propagates information on the utility function of a data point to its neighbors. Finally, they perform data selection through an iterative reverse message passing process, where at each step the highest scoring data point is kept, and all of its neighbors get down-weighted to encourage the selection of diverse data points. Lang et al. (2022) also use a graph representation as part of their utility function, determining which weakly-labeled data points to keep based on their nearest neighbors, and the label of their nearest neighbors. They only keep data points whose nearest neighbors have mostly the same label, thereby using the representations geometry to identify accurate subsets and removing ambiguous data points. Bejan et al. (2023) develop the task-agnostic self-influence score as a utility function for filtering data (and designing a curriculum). The intuition behind the self-influence score is: if not training on data point $x^{(i)}$ leads to a significantly higher loss on that data point, then there must not be many supporting data points around $x^{(i)}$ and it is likely to be an outlier, and therefore a low-quality data point. They suggest that self-influence functions can identify data with label noise, out-of-distribution data, data that is ambiguous, and difficult to learn samples, and removing such data can improve the models performance.
Many methods of improving data efficiency exist in general machine learning settings (e.g., data pruning and coreset selection), but most have exclusively been developed and evaluated in the computer vision domain. Azeemi et al. (2023) is one of the limited tests of data pruning on language tasks, where they observe that selecting for difficult to learn examples leads to improved model performance, while reducing the dataset size by up to 60%. The many other works on data pruning and coreset selection (further discussed in §9) may provide effective methods of improving data efficiency for task-specific fine-tuning, but could require some alterations to work for NLP tasks.
@@
Finally, an important problem that data selection can be applied to is improving a model’s robustness by ensuring that a model is not biased towards any particular subpopulation of data and to avoid relying on spurious correlations. Because these methods concern themselves with subgroups of data, they share similarities with data mixing (§3.8). Sagawa et al. (2020) introduce group distributionally robust optimization (group DRO) to train models that minimize the worst-case loss over groups in the training data. As their utility function, they use a notion of the generalization gap (the difference between training and validation loss) and prioritize training on data from those subgroups which have a large generalization gap. Haghtalab et al. (2022) also aim to minimize the worst-case error of a model across subgroups, while also aiming to improve the data efficiency by minimizing the number of samples required for each subgroup. They formulate the multi-distribution learning problem as a zero-sum game and aim to learn the minimax equilibrium by introducing an online learning algorithm, Resampling-based Multi-Distribution Learning (R-MDL) that uses estimates of the training error and generalization error as the utility.
@@
Compared with data selection for pretraining, the goals of data selection for task-specific fine-tuning are very similar: improving data efficiency, selecting a diverse set of samples, and reducing a model’s bias towards any particular subgroup. However, the task-specific datasets used in fine-tuning are often small enough that methods with high computational requirements are plausible. Though many of the methods presented in this section may be beneficial for pretraining (or other settings), they may be unreasonably costly to use on web-scale data. For example, Azeemi et al. (2023) requires training a model on the full dataset prior to determining which data points to keep/remove, and Haghtalab et al. (2022) assumes that the model will be validated on each subgroup at every step of the learning algorithm. Although these ideas may be computationally prohibitive to be used as-is, it may be possible to adjust them to work at larger scales by reducing the frequency of updates, or using smaller models as proxies.
@@ 8 Data Selection for Other Domains
This survey focuses on data selection for language models, but data selection is an active area of research in computer vision, vision-and-language and broadly in machine learning. Some of the data selection methods covered by this survey are domain-specific (e.g., the total number of characters in $x^{(i)}$), while others are domain agnostic (e.g., selecting data most similar to a target distribution). Therefore, some of the methods previously presented in this work may be more or less applicable in other domains. Similar to language models, data selection methods have been applied at multiple stages of model training in other domains (e.g., vision-language model pretraining) and for similar purposes (e.g., data efficiency, model performance, and reducing bias). In this section we will briefly discuss some of the prominent uses of data selection in non-language domains and compare and contrast them with the previously discussed methods.
@@
In addition to data collection and filtering, data mixing is also a topic of concern for non-language pretraining. Piergiovanni et al. (2023) propose an online data mixing method for vision-language models that mixes a variety of tasks together (e.g., image captioning or visual question answering). Their method, dynamic difficulty sampling, calculates the loss of each task $t$, $\mathcal{L}_{t}$, and weights each task proportionally with the total loss $L$, getting the final sampling weight as $S_{t}=\dfrac{\mathcal{L}_{t}}{L}$. This mixing method maintains some similarities to the online data mixing methods from language model pretraining (Xia et al., 2024a; Albalak et al., 2023a; Schioppa et al., 2023).
@@
The popularity of task-specific fine-tuning across modalities and domains has led to many data selection methods that target this setting. For the most part, the methods of data selection for non-language domains are quite similar to the methods for language. Generally speaking, these methods use the output of a model as the representation upon which the utility function is built. In an early work on support vector machines, Wang et al. (2005) devise a method of selecting data that reduces the training set size while maintaining generalization by filtering out data points which do not affect the final decision function. This idea resonates in the language domain, with similarities to methods that use loss (Albalak et al., 2023a) or methods that attempt to estimate data difficulty (Mindermann et al., 2022). A more recent direction has been how to best utilize synthetic data, and Yuan et al. (2024a) demonstrate that generating large quantities of synthetic images and selecting data using distribution matching (matching the synthetic data to the target data using maximum mean discrepancy) leads to models that perform well for image classification with synthetic data only, as well as when mixing real and synthetic data.
@@
A number of methods have been developed to improve the data selection on the level of individual batches, sometimes called data difficulty estimation. One of the motivations of these methods is that they use the current model’s state to determine which data to select, but this can add significant computational overhead. Additionally, because these methods are run alongside the model training, their cost is not amortized across training runs as it would be for static datasets. Chang et al. (2017) propose two methods for Active Bias that emphasize training on data points where the model is uncertain, based on their prediction variance or closeness to the decision threshold. Similarly, Song et al. (2020) propose Recency Bias, a method that further emphasizes training on uncertain data points that were predicted inconsistently in recent iterations of training. Finally, Mindermann et al. (2022) present Reducible Holdout Loss Selection (RHO-Loss), a method of adaptive batch selection that, intuitively, avoids selecting data points that are redundant, noisy, or less relevant. RHO-Loss was validated on a very small sample of NLP tasks, task-specific fine-tuning on SST2 (Socher et al., 2013) and CoLA (Warstadt et al., 2019), but has not been demonstrated to work in other learning settings, and Kaddour et al. (2023) show that it is not computationally efficient.
Adaptive batch selection methods are similar to online data selection, except that they are often performed on the level of individual data points, rather than groups, domains, or tasks. This particular direction has found minimal success with LLMs, in particular for pretraining (Kaddour et al., 2023), but the success of such methods in other domains suggests that it would be worthwhile adapting these methods to language.
@@
There have been a number of recent works that aim to prepare fair datasets. Biswas & Rajan (2021) analyze stage-specific fairness (SF) which measures fairness metrics with and without certain data preprocessing stages. González-Zelaya et al. (2023) introduced FairPipes, a heuristic approach to find an optimal data preparation pipeline configuration. A number of works have also developed software tools for tracing distribution distortions linked to fairness in the data preparation workflow (Yang et al., 2020a; Grafberger et al., 2021; Schelter et al., 2019).
Additionally, there exist fairness-aware balancing techniques (Yan et al., 2020; Sonoda, 2023; Chakraborty et al., 2021; Salazar et al., 2021; Yang et al., 2020b) as well as reweighting/resampling methods aimed at ensuring fairness (Kamiran & Calders, 2012; Jiang & Nachum, 2020). In NLP, using balanced training for fairness has been explored (Wang et al., 2019; Han et al., 2022). Jeong et al. (2022) showed that different proportions of demographics can lead to nontrivial differences in fairness. To alleviate the challenge of gathering additional data, recent research has focused on designing adversarial networks. These networks are developed to acquire a balanced representation, which helps in minimizing the adverse effects of sensitive attributes (Madras et al., 2018; Adel et al., 2019).
@@
Recently, the importance of data selection and quality has gained notoriety, and thus a few efforts have been developed to try and lower the barrier to entry for research to be done. One such effort is DataComp (Gadre et al., 2023), a benchmark for multimodal dataset design. DataComp provides two tracks for participation: the filtering track, and the bring-your-own-data track. The filtering track is of particular interest as it gives participants a common dataset (12.8B image-text pairs from Common Crawl), and the participants’ goal is to determine the best subset to train a model on. Additionally, Gadre et al. (2023) provide over 300 baseline experiments, and find that, similar to the language-only setting, a smaller, more selective dataset can lead to models that generalize better than larger datasets. Of course, this still comes with the same limitations as all previously discussed sections, where the exact evaluation setting plays a large role in what data is “best”. Specifically, DataComp evaluates models’ abilities to perform image classification and retrieval tasks, but does not evaluate generative capabilities. This suggests that distribution matching methods may fare well on DataComp, while an evaluation of generative capabilities may prefer data selection methods that favor distribution diversification. DataPerf (Mazumder et al., 2023) provides another entry for data selection research. The DataPerf benchmark contains 4 data selection settings: vision, speech, debugging, and language. The various settings allow researchers to test their algorithms for selection, augmentation, quality assesment, and cleaning. Mazumder et al. (2023) find that automated methods tended to perform better than manual cleaning or filtering; specifically, they found that submissions succeeded with automated methods for recognizing noisy images and labels, identifying mislabeled images, correcting class imbalance, and selecting and enhancing images from the long-tailed distribution of classes.